OpenStudy (liv1234):

A spherical balloon has a circumference of 21 cm. What is the approximate surface area of the balloon to the nearest square centimeter? A. 1,385 cm^2 B. 561 cm^2 C. 346 cm^2 D. 140 cm^2

OpenStudy (zzr0ck3r):

What is the equation for the area of a sphere?

OpenStudy (liv1234):

What do you mean?

OpenStudy (zzr0ck3r):

Area \(A\) is given by \[A=4\pi r^2\] Can you figure out the radius?

OpenStudy (liv1234):

If I knew what the diameter was I could.

OpenStudy (liv1234):

Or is there another way I could?

OpenStudy (zzr0ck3r):

\(C=2\pi r^2\)

OpenStudy (liv1234):

Right, I have the circumference which is 21. I'm trying to figure out the answer.

OpenStudy (zzr0ck3r):

So given we have \(A=4\pi (\sqrt{\frac{C}{2\pi}})^2\)

OpenStudy (liv1234):

Can you make this simpler for me please aha?

OpenStudy (zzr0ck3r):

Use the Circumfrence formula to solve for r then plug that into the other one

OpenStudy (zzr0ck3r):

\(C=2\pi r^2\) you are given \(C=21\) so we have \(21=2\pi r^2\) and we solve for \(r\) and we get \(r = \sqrt{\frac{21}{2\pi}}\) Now we plug that into the Surface area formula I showed you the first time \(A = 4\pi r^2 = 4\pi (\sqrt{\frac{21}{2\pi}})^2= 4\pi \frac{21}{2\pi}=42\)

OpenStudy (liv1234):

So, basically my radius is 42?

OpenStudy (zzr0ck3r):

no

OpenStudy (liv1234):

I hate just asking for a straight out answer, but could you just give me the answer please?

OpenStudy (zzr0ck3r):

I actually did. The fact that you dont know that is probably not a good thing...

OpenStudy (zzr0ck3r):

Which part do you now understand about what I wrote?

OpenStudy (zzr0ck3r):

now = not

OpenStudy (anonymous):

Sorry zz but circumference is C = 2*pi*r no? :P

OpenStudy (zzr0ck3r):

err yes

OpenStudy (liv1234):

I'm confused because my brain is currently going crazy thinking about my grandfather who just passed away and the fact that I really wanna finish catching back up with school since I'm 56 lessons behind..

OpenStudy (zzr0ck3r):

\(A = 4\pi r^2 = 4\pi (\frac{21}{2\pi})^2\)

OpenStudy (zzr0ck3r):

OK Liv We have \(C=2\pi r\) and we are given circumference, so we have \(21 = 2\pi r\) Are you with me so far?

OpenStudy (liv1234):

Yes.

OpenStudy (zzr0ck3r):

Ok so now we solve for \(r\) and get \(r=\frac{21}{2\pi}\) Still with me?

OpenStudy (liv1234):

Yeah, so 21/2*3.14?

OpenStudy (zzr0ck3r):

that would be an approximate. But that part is not important right now. Just know \(r=\frac{21}{2\pi}\) ok?

OpenStudy (liv1234):

Okay.

OpenStudy (zzr0ck3r):

So now remember the Surface area of a sphere is \(A= 4 \pi r^2\) right?

OpenStudy (liv1234):

Okay, that makes sense.

OpenStudy (zzr0ck3r):

Before we didnt know \(r\) so we could not solve, but now we know what \(r\) is.

OpenStudy (zzr0ck3r):

so plug in for \(r\) and we get \[A = 4 \pi (\frac{21}{2\pi})^2\]

OpenStudy (zzr0ck3r):

That is the answer

OpenStudy (liv1234):

So, it would be A?

OpenStudy (liv1234):

Wow, wait I think I put the numbers in wrong..

OpenStudy (zzr0ck3r):

what do you mean it would be A?

OpenStudy (liv1234):

Yeah, I know I messed up.

OpenStudy (zzr0ck3r):

The answer is \(4\pi (\frac{21}{2\pi})^2\approx 140.37\)

OpenStudy (liv1234):

Ohhh. Wow.